# Unscrambling of single-particle wave functions in systems localized through disorder and monitoring

### v1

@@ Abstract
In systems undergoing localization-delocalization quantum phase transitions due to disorder or monitoring, there is a crucial need for robust methods capable of distinguishing phases and uncovering their intrinsic properties. In this work, we develop a process of finding a Slater determinant representation of free-fermion wave functions that accurately characterizes localized particles, a procedure we dub “unscrambling”. The central idea is to minimize the overlap between envelopes of single-particle wave functions or, equivalently, to maximize the inverse participation ratio of each orbital. This numerically efficient methodology can differentiate between distinct types of wave functions: exponentially localized, power-law localized, and conformal critical, also revealing the underlying physics of these states. Furthermore, we apply this approach to a more challenging problem involving disordered monitored free fermions in one dimension, where the unscrambling process unveils the presence of a conformal critical phase and a localized area-law quantum Zeno phase. Importantly, our method can also be extended to free fermion systems without particle number conservation, which we demonstrate by estimating the phase diagram of $\mathbb{Z}_{2}$-symmetric disordered monitored free fermions. Our results unlock the potential of utilizing single-particle wave functions to gain valuable insights into the localization transition properties in systems such as monitored free fermions and disordered models.
@@ I Introduction
In the realm of closed quantum mechanical systems, unitary evolution often leads to thermalization, a process where information about initial conditions becomes inaccessible in local observables [1, 2, 3, 4]. There are however intriguing exceptions, where the system instead exhibits non-equilibrium behavior, such as retaining useful information for arbitrarily long times. Paradigmatic examples of this are localized systems, where the disorder is responsible for the breaking of the eigenstate thermalization hypothesis.
Within the context of many-body systems, the transition between a many-body localized (MBL) phase and a thermalized phase has captivated researchers and sparked debates [5, 6, 7, 8, 9]. The prospect of stabilizing quantum information using disorder is particularly relevant for experiments involving quantum memories, and quantum computation in general. From the analytical point of view, the transition to the localized phase can be characterized by properties of the energy spectrum and quantum states. In the former, one often compares the system to the relevant random matrix theory ensemble, examining statistics such as nearby level spacings [10, 11, 12]. On the other hand, assessing the localization properties of wave functions can be done through inverse participation ratio (IPR), a powerful tool in the study of MBL systems [13, 14, 15, 16, 17].
In the domain of free fermion systems, IPR has proven effective in characterizing the Anderson localization transition [18, 19, 20, 21], where the single-particle wave functions are exponentially localized. Albeit the single-particle wave functions serve as a valuable theoretical tool for visualizing and understanding free-fermion physics, their precise definition is somewhat ambiguous, which leaves the physics unchanged but may hamper the theoretical interpretation. As we will see in the course of this paper, we can resolve this ambiguity by establishing a clear link between the procedure for obtaining relevant single-particle representations and the IPR.
Figure 1: A diagrammatic picture of the unscrambling method $Q$, illustrating how the original Slater determinant matrix $U$ transforms into a collection of localized orbitals $U^{\prime}=UQ$.
@@
Hence, we pose the central problem of this study: is there a method of transforming the Slater determinant matrix $U$ into a set of single-particle wave functions that are endowed with specific properties, such as localization? To answer this question, we propose a methodology based on the average IPR of all particles in the system. This procedure aims to yield single-particle wave functions that reveal insights into the presence or absence of localization (as illustrated in Fig. 1), as well as internal model-specific characteristics. Furthermore, we intend to validate this method by applying it to wave functions produced through the interplay of disorder and monitoring. These systems have recently garnered attention in the context of measurement-induced entanglement transitions [22, 23, 24, 25, 26], where repeated measurements force the system into an area-law quantum Zeno phase. Our objective is to demonstrate that the proposed approach is applicable to a diverse range of systems, extending beyond those solely localized through the disorder.
The paper is structured as follows. In Sec. II we propose a methodology that approximates matrix $Q$ needed to transform the representation into localized single-particle orbitals. We test this method in Sec. III on three models exhibiting three distinct behaviors: exponentially localized, conformally invariant and delocalized, and power-law localized. In Sec. IV we show how the method behaves in a monitored disordered system with a measurement-induced phase transition between localized and delocalized regimes. We also extend the applicability of this methodology to systems without number conservation, as discussed in Sec. V. Finally, we conclude in Sec. VI.
@@ II.1 The two-particle problem
Now, we introduce a cost function that encompasses the localization properties of the two particles. We propose minimizing the overlap between the envelopes of single-particle wave functions. For two particles labeled 1 and 2, this cost function is defines as
@@ II.2 The multi-particle problem
where IPR of a particle is defined as $\text{IPR}(n)=\sum_{i}|U_{in}|^{4}=\sum_{i}|\langle i|\psi_{n}\rangle|^{4}$. IPR serves as a measure of localization [18, 19, 20, 21], and can take values between 0 and 1. When IPR is large, the orbital is localized in the given basis (in the real space, $\text{IPR}\sim 1/\xi\sim\text{const.}$. for localization length $\xi$), while if IPR is approximately $1/L$, the orbital is delocalized, suggesting thermalization. This gives the physical interpretation of the unscrambling method: by minimizing the global cost function $f$, which represents sums of overlaps between single-particle wave function envelopes, we maximize the average IPR [defined as $\sum_{n}\text{IPR}(n)/N$] of the particle orbitals. Intuitively, this approach ensures that the orbitals become as localized as possible.
@@ III.1 Exponentially localized wave functions in Anderson localized systems
$U(t+dt)=e^{i\,dt\,\mathbb{H}}U(t),$ (19)
where $\mathbb{H}$ is an $L\times L$ matrix representing the Hamiltonian, with elements $\mathbb{H}_{ij}=\delta_{i,j+1}+\delta_{i,j-1}+h_{i}\delta_{ij}$. We set the time step to $dt=0.05$. At long times, this evolution should lead to a wave function, which can be represented as a Slater determinant of exponentially localized single-particle wave functions. We, therefore, evolve the system up to $t_{\text{fin}}=100$ and subsequently apply the unscrambling method.
Figure 3: Anderson localization of single-particle wave functions revealed after the unscrambling procedure. The dashed lines is the expected behavior of $\sim\exp(-|i|/\xi)$, with $\xi\approx 25/W^{2}$. $L=204,dt=0.05,t_{\text{fin}}=100$.
@@
where $\bar{\ast}$ is an average over 100 disorder realizations and all particles $n$.
The resulting orbitals exhibit the expected exponential localization, $\overline{\overline{|\psi_{i}|^{2}}}\sim\exp(-|i|/\xi)$. The behavior of the localization length $\xi\approx 25/W^{2}$ is in complete agreement with the well-known behavior (see the dashed lines in the figure) for Anderson localization, including the prefactor [27, 18]. This simple test shows that the unscrambling procedure works well for exponentially-localized wave functions, and correctly picks up the value of the localization length.
@@ III.2 Conformally-invariant wave functions in the XX model
We now investigate the properties of the unscrambled single-particle wave functions. To obtain $U$ from $D$ one can use singular value decomposition, where the singular values should be either 0 or 1. Fig. 5 shows the orbitals for a few values of the transverse field $h$ and a set system size $L=204$. The number of occupied modes is $N=89$ for $h=0.4$, $N=47$ for $h=1.5$, and $N=21$ for $h=1.9$. We find two interesting features of the orbitals. First, the single-particle wave functions show oscillations with the number of troughs equal to $(N-1)$. This feature seems to be tightly related to the specifics of the model itself. We leave for future research the peculiar features of these oscillations, such as why each oscillation has a dome-like structure.
Second, the orbitals on average follow a $\overline{\overline{|\psi_{i}|^{2}}}\sim\csc^{2}(i\pi/L)$ behavior. This can be related to the correlation functions from Eq. (25), which are modulated by $\sim\csc[(i-j)\pi/L]$. In the infinite-volume limit, this turns into a $\sim 1/(i-j)$ behavior. This brings us to one of properties of the critical phase often used as its signature: the connected correlation functions, defined as $C(r)=|D_{i,i+r}|^{2}=\langle n_{i}\rangle\langle n_{i+r}\rangle-\langle n_{i}n_{i+r}\rangle$, should decay algebraically as $\sim 1/r^{2}$ in the conformal phase.
@@ III.3 Power-law localization in the symplectic Hatano-Nelson model
We plot the spin-up and spin-down parts of the unscrambled orbitals in Fig. 6. Each wave function either faces right or left, i.e. its magnitude grows towards right or towards left of the system. We can also see that there is a difference between odd and even sites of the chain: one sublattice faces right/left, while the other is either close to zero or constant. This shows an imbalance between the right/left-facing modes, which is the origin of the skin effect. We also show power-law localization by fitting an ansatz $\overline{\overline{|\psi_{i}|}}\sim a-b|i|^{\alpha}$, where $a,b,\alpha$ are fitting parameters, and $\alpha$ is the power-law exponent. The fitting is done only to the sublattice that is not constant. The fitted power law exponent is consistently found to be $\alpha\approx 0.88$ for all wave functions.
In summary, the unscrambling method yields several significant insights into the underlying properties of the wave functions of the symplectic Hatano-Nelson model. First, unscrambling reveals the existence of two types of quasiparticles – one traveling right, the other left. Second, it clearly shows the difference between the even and odd sublattices of the system. Third, the unscrambled orbitals are indeed power-law localized.
IV Single-particle Wave functions in disordered monitored free fermions
@@
where $\gamma$ is the measurement strength, $d\eta_{i}^{t}$ is an Itô increment (a random number with variance ${\gamma\,dt}$), and the Hamiltonian $H$ is that of the Anderson model from Eq. (21). This stochastic evolution models a continuous monitoring of particle number occupation, which can be imagined as a series of homodyne detectors in an experiment [35]. The dynamics can be translated into the following change in the matrix $U$ of single-particle wave functions:
@@
We initialize the wave function as a half-filled Néel state and stop the evolution after the long-time steady state is reached, $t_{\text{fin}}=8L$ for $W\leq 1$, and $t_{\text{fin}}=8LW$ when $W>1$. We subsequently unscramble the resulting wave functions, and plot them in Fig. 7. Examples of wave functions with parameters chosen inside the critical phase are in Fig. 7(a), and all fit well within an ansatz $\overline{\overline{|\psi_{i}|^{2}}}\sim\csc^{2}(i\pi/L)$ (dashed lines). This shows that the wave functions exhibit the generic properties of criticality that we discussed in Sec. III.2. The fit does not work well only near the edges of the plot, which correspond to short distances from the center of the wave function, where we indeed do not expect generic behavior. Furthermore, Fig. 7(b) shows parameters within the area-law phase, where the single-particle wave functions exhibit exponential localization, as in Sec. III.1. Fitting an exponential decay (dashed lines) fails near the (localization) center, but also near the very middle of the plot, where large distances are impacted by finite-size effects related to periodic boundary conditions. A similar behavior was seen in Sec. III.1.
@@
V Unscrambling without particle number conservation

### v3

@@ Abstract
In systems undergoing localization-delocalization quantum phase transitions due to disorder or monitoring, there is a crucial need for robust methods capable of distinguishing phases and uncovering their intrinsic properties. In this work, we develop a process of finding a Slater determinant representation of free-fermion wave functions that accurately characterizes localized particles, a procedure we dub “unscrambling”. The central idea is to minimize the overlap between envelopes of single-particle wave functions or, equivalently, to maximize the inverse participation ratio of each orbital. This numerically efficient methodology can differentiate between distinct types of wave functions: exponentially localized, power-law localized, and conformal critical, also revealing the underlying physics of these states. The method is readily extendable to systems in higher dimensions. Furthermore, we apply this approach to a more challenging problem involving disordered monitored free fermions in one dimension, where the unscrambling process unveils the presence of a conformal critical phase and a localized area-law quantum Zeno phase. Importantly, our method can also be extended to free fermion systems without particle number conservation, which we demonstrate by estimating the phase diagram of $\mathbb{Z}_{2}$-symmetric disordered monitored free fermions. Our results unlock the potential of utilizing single-particle wave functions to gain valuable insights into the localization transition properties in systems such as monitored free fermions and disordered models.
@@ I Introduction
In closed quantum mechanical systems, unitary evolution often leads to thermalization, a process where information about initial conditions becomes inaccessible in local observables [1, 2, 3, 4]. There are however intriguing exceptions, where the system instead exhibits non-equilibrium behavior, such as retaining useful information for arbitrarily long times. Paradigmatic examples of this are localized systems, where the disorder is responsible for the breaking of the eigenstate thermalization hypothesis.
Within the context of many-body systems, the transition between a many-body localized (MBL) phase and a thermalized phase has captivated researchers and sparked debates [5, 6, 7, 8, 9]. The prospect of stabilizing quantum information using disorder is particularly relevant for experiments involving quantum memories, and quantum computation in general. From the conceptual point of view, the transition to the localized phase can be characterized by properties of the energy spectrum and quantum states. In the former, one often compares the system to the relevant random matrix theory ensemble, examining statistics such as nearby level spacings [10, 11, 12]. On the other hand, assessing the localization properties of wave functions can be done through inverse participation ratio (IPR), a powerful tool in the study of MBL systems [13, 14, 15, 16, 17].
In the case of free fermion systems, IPR has proven effective in characterizing the Anderson localization transition [18, 19, 20, 21], where the single-particle wave functions become exponentially localized in the insulating phase. Other non-equilibrium processes in free fermion systems have seen a resurgence in interest, owing largely to their tractability through the single-particle picture. Most prominent examples include dynamical phase transitions [22, 23, 24, 25], time crystals [26, 27], quantum quenches [28, 29, 30, 31, 32, 33, 34], Floquet engineering [35, 36, 37, 38, 39, 40], and monitored quantum systems [41, 42]. The last have garnered attention in the context of measurement-induced entanglement transitions [43, 44, 45, 46, 47, 48], where repeated measurements force the system into an area-law quantum Zeno phase. Notably, albeit the single-particle wave functions serve as a valuable theoretical tool for visualizing and understanding free-fermion physics, their precise definition is somewhat ambiguous, which leaves the physics unchanged but may hamper the theoretical interpretation. As we will see in the course of this paper, we can resolve this ambiguity by establishing a clear link between the procedure for obtaining relevant single-particle representations and the IPR.
Figure 1: A diagrammatic picture of the unscrambling method $Q$, illustrating how the original Slater determinant matrix $U$ transforms into a collection of localized orbitals $U^{\prime}=UQ$. $U$ and $U^{\prime}$ are represented by plotting the density of the corresponding orbitals as a function of a position. $Q$ is approximated as a series of two-orbital transformations (green gates).
@@
Hence, we pose the central problem of this study: is there a method of transforming the Slater determinant matrix $U$ into a set of single-particle wave functions that are endowed with specific properties, such as localization? To answer this question, we propose a methodology based on the average IPR of all particle orbitals in the system. This procedure aims to yield single-particle wave functions that reveal insights into the presence or absence of localization (as illustrated in Fig. 1), as well as internal model-specific characteristics. Furthermore, we intend to validate this method by applying it to wave functions produced through the interplay of disorder and monitoring, where a measurement-induced entanglement transition is present. Our objective is to demonstrate that the proposed approach is applicable to a diverse range of systems, extending beyond those solely localized through the disorder.
The paper is structured as follows. In Sec. II we propose a methodology that approximates matrix $Q$ needed to transform the representation into localized single-particle orbitals. We test this method in Sec. III on three models exhibiting three distinct behaviors: exponentially localized, conformally invariant, and power-law localized. We apply the method to a higher-dimensional system in Sec. IV, showing its versatility across different lattice geometries. In Sec. V we show how the method behaves in a monitored disordered system with a measurement-induced phase transition between localized and critical regimes. We also extend the applicability of this methodology to systems without number conservation, as discussed in Sec. VI. Finally, we conclude in Sec. VII.
@@ II.1 The two-particle problem
Now, we introduce a cost function that encompasses the localization properties of the two particles. We propose minimizing the overlap between the envelopes of single-particle wave functions. For two particles labeled 1 and 2, this cost function is defined as
@@ II.2 The multi-particle problem
where IPR of a single-particle wave function is defined as $\text{IPR}(n)=\sum_{i}|U_{in}|^{4}=\sum_{i}|\langle i|\psi_{n}\rangle|^{4}$. IPR serves as a measure of localization [18, 19, 20, 21], and can take values between 0 and 1. When IPR is large, the orbital is localized in the given basis (in the real space, $\text{IPR}\sim 1/\xi\sim\text{const.}$. for localization length $\xi$), while if IPR is approximately $1/L$, the orbital is delocalized (which is a necessary, but not a sufficient condition for thermalization). This gives the physical interpretation of the unscrambling method: by minimizing the global cost function $f$, which represents sums of overlaps between single-particle wave function envelopes, we maximize the average IPR [defined as $\sum_{n}\text{IPR}(n)/N$] of the particle orbitals. Intuitively, this approach ensures that the orbitals become as localized as possible.
@@ III.1 Exponentially localized wave functions in Anderson localized systems
$U(t+dt)=e^{-i\,dt\,\mathbb{H}}U(t),$ (19)
where $\mathbb{H}$ is an $L\times L$ matrix representing the Hamiltonian, with elements $\mathbb{H}_{ij}=\delta_{i,j+1}+\delta_{i,j-1}+h_{i}\delta_{ij}$. We set the time step to $dt=0.05$. At long times, this evolution should lead to a wave function, which can be represented as a Slater determinant of exponentially localized single-particle wave functions. We, therefore, evolve the system up to $t_{\text{fin}}=100$ and subsequently apply the unscrambling method. Note that alternatively, we could use the correlation matrix $D$ and evolve it as $D(t+dt)=\exp(-i\,dt\,\mathbb{H})D\exp(i\,dt\,\mathbb{H})$, then subsequently use singular value decomposition to obtain $U$. However, applying operations directly on $U$ is numerically much more efficient.
Figure 3: Anderson localization of single-particle wave functions revealed after the unscrambling procedure. The dashed lines is the expected behavior of $\sim\exp(-|i|/\xi)$, with $\xi\approx 20/W^{2}$. $L=204,dt=0.05,t_{\text{fin}}=100$.
@@
where $\bar{\ast}$ is an average over 100 disorder realizations and all particle orbitals $n$. This method of averaging (i.e. centering first, then taking the typical average) will be used throughout the rest of this paper.
The resulting orbitals exhibit the expected exponential localization, $\overline{\overline{|\psi_{i}|^{2}}}\sim\exp(-|i|/\xi)$. The behavior of the localization length $\xi\approx 20/W^{2}$ is in complete agreement with the well-known behavior (see the dashed lines in the figure) for Anderson localization, including the prefactor [49, 18]. This simple test shows that the unscrambling procedure works well for exponentially-localized wave functions, not only revealing localization, but also correctly predicting the value of the localization length.
@@ III.2 Conformally-invariant wave functions in the XX model
We now investigate the properties of the unscrambled single-particle wave functions. To obtain $U$ from $D$ one can use singular value decomposition, where the singular values should be either 0 or 1. Fig. 5 shows the orbitals for a few values of the transverse field $h$ and a set system size $L=204$. The number of occupied modes is $N=89$ for $h=0.4$, $N=47$ for $h=1.5$, and $N=21$ for $h=1.9$. We find two interesting features of the orbitals. First, the single-particle wave functions show oscillations with the number of troughs equal to $(N-1)$, which seems to be tightly related to the specifics of the model itself. We leave for future research the peculiar features of these oscillations, such as why each oscillation has a dome-like structure.
Second, the orbitals on average follow a $\overline{\overline{|\psi_{i}|^{2}}}\sim\csc^{2}(i\pi/L)$ behavior. This can be related to the correlation functions from Eq. (25), which are modulated by $\sim\csc[(i-j)\pi/L]$. In the infinite-volume limit, this turns into a $\sim 1/(i-j)$ behavior. This brings us to one of properties of the critical phase often used as its signature: the connected correlation functions, defined as $C(r)=|D_{i,i+r}|^{2}=\langle n_{i}\rangle\langle n_{i+r}\rangle-\langle n_{i}n_{i+r}\rangle$, should decay algebraically as $\sim 1/r^{2}$ in the conformal phase. Note that although the unscrambled wave functions are critical, the single-particle eigenstates of this model will be delocalized.
@@ III.3 Power-law localization in the symplectic Hatano-Nelson model
Figure 7: Unscrambled single-particle wave functions of the 3D Anderson model, with the disorder strength of (a) $W=4$ (diffusive) and (b) $W=12$ (localized). We use a lattice of $10\times 10\times 10$ sites. The legend in panel (b) applies in panel (a). (c) Unscrambled orbital dependence on the distance $r=\sqrt{x^{2}+y^{2}+z^{2}}$ from the localization center. The dashed line shows the exponential decay $\overline{\overline{|\psi_{i}|^{2}}}\sim\exp(-r/\xi)$, where $\xi=3.457$ [55] is the localization length of the infinite system for $W=12$.
@@
We plot the spin-up and spin-down parts of the unscrambled orbitals in Fig. 6. Each wave function either faces right or left, i.e. its magnitude grows towards the right or towards the left of the system. We can also see that there is a difference between odd and even sites of the chain: one sublattice faces right/left, while the other is either close to zero or constant. This shows an imbalance between the right/left-facing modes, which is the origin of the skin effect. We also show power-law localization by fitting an ansatz $\overline{\overline{|\psi_{i}|}}\sim a-b|i|^{\alpha}$, where $a,b,\alpha$ are fitting parameters, and $\alpha$ is the power-law exponent. The fitting is done only to the sublattice that is not constant. The fitted power law exponent is consistently found to be $\alpha\approx 0.88$ for all wave functions.
In summary, the unscrambling method yields several significant insights into the underlying properties of the wave functions of the symplectic Hatano-Nelson model. First, unscrambling reveals the existence of two types of quasiparticles – one traveling right, the other left. Second, it clearly shows the difference between the even and odd sublattices of the system. Third, the unscrambled orbitals are indeed power-law localized, as predicted by the theory.
IV Unscrambling in higher dimensions
One interesting advantage of the unscrambling methodology is that the cost function is essentially independent of the lattice geometry. Consequently, the procedure can be readily extended to higher dimensional lattices. In this section, we apply the unscrambling method to the three-dimensional (3D) Anderson model [18, 56, 57]. In contrast to the one-dimensional case from the previous section, as well as the two-dimensional case, this model exhibits an Anderson phase transition from a diffusive metallic phase to a localized insulating phase.
We formally define the 3D Anderson model by the Hamiltonian from Eq. (18), with the hopping term applied between each link on a cubic lattice of size $L\times L\times L$, and with periodic boundary conditions. The Anderson transition happens at the critical disorder strength of $W_{c}\approx 8.27$ [58]. We initialize the wave function to be half-filled in a 3D checkerboard pattern and evolve the system for time $t_{\text{fin}}=8L$. Subsequently, we perform the unscrambling.
The unscrambled orbitals are shown in Fig. 7. The averaging is done similarly to the one-dimensional case, where we center each orbital at position $(0,0,0)$ before averaging. The unscrambled single-particle wave functions show a clear change between the diffusive regime [$W=4$, panel (a)], where the orbitals are extended across the entire lattice, and the localized phase [$W=12$, panel (b)], where the orbitals are localized. In Fig. 7(c), we show the orbitals as a function of the distance from the localization center, $r=\sqrt{x^{2}+y^{2}+z^{2}}$. In the diffusive phase, the orbital is essentially constant, except for $r=0$, which is an artifact of the averaging procedure. On the other hand, in the localized phase, the orbital exhibits exponential decay $\overline{\overline{|\psi_{i}|^{2}}}\sim\exp(-r/\xi)$, where the localization length $\xi=3.457$ for $W=12$ is taken from the literature [55]. The fit works best away from the localization center, where the behavior is expected to be model-dependent, and away from the maximal distances, where finite-size effects are expected to dominate.
Summarizing, the unscrambling methodology is shown to work in the three-dimensional Anderson model, correctly identifying the presence of the diffusive and localized phases, and predicting the wave function decay consistent with the literature.
V Single-particle Wave functions in disordered monitored free fermions
@@
where $\gamma$ is the measurement strength, $d\eta_{i}^{t}$ is an Itô increment (a random number with variance ${\gamma\,dt}$), and the Hamiltonian $H$ is that of the Anderson model from Eq. (18). This stochastic evolution models a continuous monitoring of particle number occupation, which can be imagined as a series of homodyne detectors in an experiment [42]. The dynamics can be translated into the following change in the matrix $U$:
@@
We initialize the system as a half-filled Néel state and stop the evolution after the long-time steady state is reached, $t_{\text{fin}}=8L$ for $W\leq 1$, and $t_{\text{fin}}=8LW$ when $W>1$. We subsequently unscramble the resulting wave functions, and plot them in Fig. 8. Examples of wave functions with parameters chosen inside the critical phase are in Fig. 8(a), and all fit well within an ansatz $\overline{\overline{|\psi_{i}|^{2}}}\sim\csc^{2}(i\pi/L)$ (dashed lines). This shows that the wave functions exhibit the generic properties of criticality that we discussed in Sec. III.2. The fit does not work well only near the edges of the plot, which correspond to short distances from the center of the wave function, where we indeed do not expect generic behavior. Furthermore, Fig. 8(b) shows parameters within the area-law phase, where the single-particle wave functions exhibit exponential localization, as in Sec. III.1. Fitting an exponential decay (dashed lines) fails near the (localization) center, but also near the very middle of the plot, where large distances are impacted by finite-size effects related to periodic boundary conditions. A similar behavior was seen in Sec. III.1.
@@
VI Unscrambling without particle number conservation